OpenStudy (anonymous):

Equation of the tangent line at the indicated point (x+1)^2 + (y-2)^2 = 20 point(3,4)

OpenStudy (gorv):

find the derivative with respect to x...which wiill give you dy/dx

OpenStudy (gorv):

that is slope of the line and plug (3,4) to find slope at this point now u have slope of tangent

OpenStudy (gorv):

and one point ....u can use point slope form of line y=mx+c

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

I getting confused with y does : y dy/dx = 2(x+1)+2(y-2) (1) + (y dy/dx) = 0, then I start plugging ???